OpenStudy (anonymous):

what is the y-intercept of the graph of the given line by the equation 3y+4x=15 A;5 B; 4 C: 3 D: -4 over 3 @Rhystic

Nnesha (nnesha):

y = mx + b this is slope intercept form where m is slope which is always with x and b is y-intercept so you have to solve for y

OpenStudy (rhystic):

so its linear but you need to solve out for y first. 3y = 15- 4x -> divide by three and then you have your y intercept

OpenStudy (anonymous):

ok so divide -15 by 3 ?

OpenStudy (rhystic):

15 by 3 yes. its not negative the 4x becomes negative because you had to bring it over to the right side.

OpenStudy (anonymous):

ok so 3 divided by 3 is 5 so -4x=5?

OpenStudy (rhystic):

no no so you have 3y = 15 - 4x you want to solve for y so you divide everything by three leaving you with y = 15/3 - 4x/3. remember that the equation of a line is y = mx + b so your y intercept is the number with neither y or x which is 15/3 so your answer is what?

OpenStudy (anonymous):

5?

OpenStudy (rhystic):

5 is correct. but -4x = 5 is not. your y intercept is 5 so the equation becomes y = 5 - 4x/3 and b = 5 your y - intercept.

OpenStudy (anonymous):

ok got it y-intercept thanks for the help @Rhystic @Nnesha